A man start from a point A and walk a distance of 8.0 km due northeast to point B from B he then walks 4.0 km due South East to point x. Calculate the shortest distance between A and x.






1
Expert's answer
2021-06-18T11:21:56-0400

"\vec{AX}=\vec{AB}+\vec{BX}"

"\angle(\vec{AB},\vec{BX})=\angle(\vec{NE},\vec{SE})=\frac{\pi}{2}"

"\vec{AX}^2=(\vec{AB}+\vec{BX})^2"

"\vec{AX}^2=\vec{AB}^2+\vec{BX}^2+2\vec{AB}*\vec{BX}"

"2\vec{AB}*\vec{BX}=2* AB*BX*\cos(\angle(\vec{AB},\vec{BX}))=0"

"AX^2 =AB^2+BX^2"

"AX =\sqrt{AB^2+BX^2}=\sqrt{8^2+4^2}=8.94"

"\text{Answer : }AX =8.94\ km"